OpenStudy (anonymous):

A line has a slope of –4 and passes through the point (8, –7). What is the equation of the line? 4x + y = 25 4x + y = 1 –4x + y = –15 –4x + y = –39 A or D?

OpenStudy (studygurl14):

slope of -4. y = mx + b y= (-4)x + b Point (8,-7) Plug in and solve for b (-7)=-4(8)+b b = 25 y = -4x + 25 4x + y = 25

OpenStudy (anonymous):

So, A?

OpenStudy (studygurl14):

yep